Exercise 123 On January 2, 2018, Tylor Company issued a 4-year, $650,000 note at 8% fixed interest, interest payable semiannually. Tylor now wants to change the note to a variable rate note. As a result, on January 2, 2018, Tylor Company enters into an interest rate swap where it agrees to receive 8% fixed and pay LIBOR of 5.7% for the first 6 months on $650,000. At each 6-month period, the variable interest rate will be reset. The variable rate is reset to 6.5% on June 30, 2018. Compute the net interest expense to be reported for this note and related swap transaction as of June 30, 2018. Net interest expense $ LINK TO TEXT Compute the net interest expense to be reported for this note and related swap transaction as of December 31, 2018. Net interest expense $
Answer #1

B 30-06-2018 650000 4% 31-12-2018 650000 4% A 1 Particulars 2 Fixed-rate debt 3 Fixed rate Semi Annual Debt 4 Payment 5 Swap1 Particulars 2 Fixed-rate debt 3 Fixed rate 650000 =8%/2 650000 =8%/2 Semi Annual Debt Payment =B2*B3 =B4 =C2*C3 =C4 10 5 Sw